# Into the Rabbit Hull: From Task-Relevant Concepts in DINO to Minkowski Geometry

### v1

@@ 1 Introduction
Trained in a self-supervised manner on a massive corpus of unlabeled images, DINOv2 has become a showcase of emergent visual representations Oquab et al. (2023); Darcet et al. (2023). Its embeddings support diverse downstream tasks: fine-grained classification Chiu et al. (2024), segmentation Liu et al. (2023), monocular depth estimation Mao et al. (2024); Cui et al. (2024), object tracking Faber et al. (2024); Tumanyan et al. (2024), even robotic perception Kim et al. (2024). They have also proven valuable in non-discriminative settings, acting as priors for generative models Yu et al. (2025), proxies for distributional similarity Stein et al. (2023), and tools for revealing blindspots Bohacek et al. (2025). The embeddings are not just visually rich, but robust: they power video world models that predict physical dynamics from frames Baldassarre et al. (2025), and can be aligned with language to achieve open-vocabulary segmentation and zero-shot classification despite having never been trained on text Jose et al. (2025). DINOv2 also transfers robustly across domains Zhang et al. (2023); Zhou et al. (2024), including satellite imagery Waldmann et al. (2025) and medical scans Ayzenberg et al. (2024), underscoring that the model has internalized a deeply structured representation of the visual world. Yet despite these successes, the nature of DINOv2’s internal representations remains unclear. What, precisely, is encoded? Which features are available Hermann et al. (2023) to downstream tasks, and how are they geometrically organized? More fundamentally, how can a label-free objective give rise to such effective few-shot readout?
@@
To address these questions, we draw on vision explainability, a field that has developed both an empirical toolkit and theoretical accounts for probing large vision models Doshi-Velez & Kim (2017); Gilpin et al. (2018); Fel (2024). Early work emphasized attribution-based methods Simonyan et al. (2013); Zeiler & Fergus (2014); Bach et al. (2015); Springenberg et al. (2014); Smilkov et al. (2017); Sundararajan et al. (2017); Selvaraju et al. (2017); Fong & Vedaldi (2017); Fel et al. (2021); Novello et al. (2022); Muzellec et al. (2024), which visualize “where” a model looks or how sensitive it is to local perturbations. While often compelling, such techniques provide only a surface-level view: they do not reveal “what” internal features are being computed, and their usefulness is limited Hase & Bansal (2020); Hsieh et al. (2021); Nguyen et al. (2021); Colin et al. (2021); Kim et al. (2022); Sixt et al. (2020). To move beyond attribution, the community turned to concept-based explanations Kim et al. (2018); Poeta et al. (2023), which aim to extract interpretable latent dimensions from a model’s representation space Bau et al. (2017); Ghorbani et al. (2019); Zhang et al. (2021); Fel et al. (2023c); Graziani et al. (2023); Vielhaben et al. (2023); Kowal et al. (2024a, b); Fel et al. (2023b). A natural first attempt was to identify individual neurons as carriers of concepts. However, this view quickly ran into challenges: neurons are polysemantic, depend on arbitrary bases, and implicitly assume the number of features is bounded by the number of units Arora et al. (2018); Elhage et al. (2022); Gorton (2024). These limitations motivated methods that identify concepts as distributed features, without requiring alignment to single neurons Ghorbani et al. (2017); Zhang et al. (2021); Fel et al. (2023c). A first theoretical account soon followed in the form of phenomenology by Elhage et al. (2022): in a $d$-dimensional space, neural networks can encode exponentially many nearly orthogonal features by representing each as a sparse combination of neurons. This phenomenology crystallized into the Linear Representation Hypothesis (LRH), which holds that models contain many more features than neurons, arranged as sparse, quasi-orthogonal directions Park et al. (2024). If the LRH is valid, then concept extraction amounts to an overcomplete dictionary learning problem Fel et al. (2023b). In this framework, the activation space of a layer is factorized into a dictionary basis and sparse codes. Sparse autoencoders (SAEs) Makhzani & Frey (2014); Cunningham et al. (2023); Bricken et al. (2023); Joseph et al. (2025a, b) are one practical instantiation of this idea, enforcing sparsity through a simple encoder-decoder architecture. When applied to pretrained models such as DINOv2, SAEs uncover a rich library of patterns, which we refer to as concepts.
@@ 2 Linear Representation Hypothesis and Operationalization
A recurring phenomenology of large models is that their representational capacity vastly exceeds the number of neurons: in a $d$-dimensional space, they can encode exponentially many features by representing each as a sparse linear superposition Arora et al. (2018); Elhage et al. (2022). Empirically, such features behave as nearly orthogonal directions Papyan et al. (2020), active only in restricted contexts, while neurons themselves are polysemantic Nguyen et al. (2016). This motivates two conclusions: (i) neurons are not the appropriate locus of interpretability, and (ii) one must recover the latent basis along which the model effectively operates. Beyond empirical characterization, such geometry can be motivated by compression: when features are arranged with minimal coherence, it maximizes the number of retrievable features while minimizing destructive interference. This principle extends beyond neural activations to fundamental problems in discrete geometry and optimization theory. It connects to classical sphere packing problems: Tammes’s problem Mooers (1994), which seeks optimal angular separation of points on a sphere (originally motivated by pollen grain morphology, where surface protrusions must be optimally spaced for aerodynamic dispersal efficiency); Thomson’s problem Thomson (1904); Bowick et al. (2002), which minimizes Coulomb electrostatic repulsion energy between charged particles constrained to a sphere; and the spherical code problem Delsarte et al. (1991), which maximizes minimum distance between codewords on the unit sphere for error correction. These problems share a common mathematical substrate: the geometric structure formalized as Grassmannian frames in signal processing theory Strohmer & Heath Jr (2003). This object underlies the optimization principle of minimizing mutual coherence (the maximum absolute inner product between distinct normalized vectors). These mathematical convergences suggest that neural networks may naturally approximate such optimal geometric configurations to (i) maximize representational capacity while (ii) minimizing cross-feature interference. This geometric phenomenology crystallizes into the Linear Representation Hypothesis Elhage et al. (2022); Costa et al. (2025) that we recall here:
@@
A persistent challenge, however, is stability: naïve SAEs produce inconsistent features across runs, undermining interpretability Paulo & Belrose (2025); Papadimitriou et al. (2025). To address this, we adopt a stable SAEs Fel et al. (2025), which constrain each dictionary atom to lie in the convex hull of real activations. This guarantees that atoms remain in-distribution and yields reproducible, geometrically faithful dictionaries. Formally, let $(\mathcal{X},\mathcal{F},\mathbb{P})$ denote the probability space of natural images, $\mathcal{X}\subset\mathbb{R}^{H\times W\times 3}$. For a pretrained Vision Transformer $\bm{f}:\mathcal{X}\to\mathbb{R}^{t\times d}$, any image $\bm{x}\sim\mathbb{P}$ yields activations $\bm{a}=\bm{f}(\bm{x})\in\mathbb{R}^{t\times d}$, i.e. $t$ token embeddings of dimension $d$. For a batch $\bm{X}=(\bm{x}_{i})_{i=1}^{n}$, concatenating all tokens gives $\bm{A}\in\mathbb{R}^{nt\times d}$. Our objective is to factorize $\bm{A}$ into sparse codes $\bm{Z}\in\mathbb{R}^{nt\times c}$ and a dictionary $\bm{D}\in\mathbb{R}^{c\times d}$ with
@@ 5 Statistics and Geometry of Concepts
However, superimposed on this sparse backdrop are three outlier concepts that exhibit an anomalously dense activation pattern Sun et al. (2025), firing across the entire dataset. Upon inspection, these concepts are not arbitrary artifacts, but instead they encode positional information: specifically indicating whether a token lies on the left, right, or bottom of the input image grid. Their ubiquity reflects the persistent relevance of spatial structure within the model, and their emergence aligns with our decoding experiments, which reveal that positional information remains linearly decodable up to the intermediate layers (see Appendix˜I) and form a 2-dimensional subspace at the penultimate layer (see Section˜6). This finding adds nuance to the assumption that the concept space is strictly sparse. Instead, we observe a hybrid representational regime where a small set of universally-active, (here spatially grounded features) coexists with a broader ensemble of highly selective, image-contingent activations, resembling sparse and dense low-rank representations Jiang & Lai (2014); Pramanik et al. (2020).
@@
Having examined the statistics of activations, we now turn to the geometry of the learned dictionary.Figure˜10 summarizes structural properties of the concept basis $\bm{D}$, including pairwise dot products, singular value spectrum, and sparsity scores. Local Coherence. Panel (A) reveals that relative to the isotropic random vectors (e.g. at initialization), the dictionary does not move towards the uniform incoherence of an ideal Grassmannian frame. Instead, the SAE dictionary moves to the other direction and exhibits higher coherence: while most atoms are weakly correlated, a non-negligible fraction form tight clusters with stronger alignment. This geometry departs from the assumptions of the LRH. Rather than approximating a Johnson–Lindenstrauss like embedding Larsen & Nelson (2017), [^4] The Johnson–Lindenstrauss (JL) lemma Johnson et al. (1984) implies that $p=\exp(\mathcal{O}(m\varepsilon^{2}))$ vectors can be embedded into $\mathbb{R}^{m}$ such that all pairwise inner products are preserved up to a small distortion $\varepsilon>0$. This supports the feasibility of near-orthogonality even in regimes where $p\gg m$. the model appears to favor structured redundancy, potentially to support specialization or compositional reuse. Moreover, a pattern of interest emerges upon inspecting the tail of the dot-product distribution: a small but significant number of concept pairs are nearly antipodal, i.e., $\bm{D}_{i}\approx-\bm{D}_{j}$ (see Figure˜11, A). This go against the assumption of a Grassmannian-like dictionary  Strohmer & Heath Jr (2003). Instead, it suggests that DINO sometimes exploits paired directions to encode semantically opposed features such as “left vs right” or “white vs black” along shared lines but with opposite signs. Practically, such antipodal structure should be accounted when using cosine similarity in downstream analyses, particularly in data filtering, clustering, or retrieval contexts. Global Anisotropy. The singular value spectrum in Panel (B) shows a sharp decay, indicating that the dictionary spans a low-dimensional subspace despite its overcomplete size. This again suggests an anisotropic allocation of representational capacity: some directions are densely populated, while others are sparsely or not at all represented. One potential explanation is that the activations $\bm{A}$ themselves have anisotropic covariance, thus to reconstruct them $\bm{A}\approx\bm{Z}\bm{D}$ with L2 error, more atoms are allocated to the more important and higher variance subspace. [^5] A speculative hypothesis is that the affine components of the last LayerNorm deforms the activation space into an ellipsoidal geometry, locally increasing curvature along some directions. In such regions, dictionary learning may require more atoms to adequately cover the manifold with low reconstruction error. While this remains to be formalized, similar intuitions appear in manifold tiling and sparse coding contexts Peyré (2009). Distributed Encoding. Finally, Panels (C) and (D) show that the dictionary atoms are not aligned to individual units: their Hoyer scores lie far below the one-hot bound. Compared to random baselines, SAE atoms are slightly sparser but remain broadly distributed—supporting the view that concept representations are shared across the population rather than axis-aligned Elhage et al. (2022); Colin et al. (2024).
@@ 6 The Shape of an Image
A natural hypothesis is that the local connectedness observed in PCA maps, illustrated in Figure˜13 and Figure˜12, is inherited from positional encodings. After all, spatial position is the only structured signal available at initialization, and remains entangled with token identity throughout the network. To test this, we begin by analyzing the evolution of positional information across layers. We trained linear decoders to decode the token position from each layer’s activations and extracted the corresponding decoding vectors at each spatial location $(i,j)$ to form a position encoding matrix $\bm{P}\in\mathbb{R}^{256\times d}$ at each layer (see Appendix˜I for more details). In early layers, this positional embedding matrix $\bm{P}$ is effectively high-rank: we can linearly decode the precise $(i,j)$ location of a token with near-perfect accuracy (see Fig. 14, Top). This aligns with the model’s need for precise positional resolution during early attention operations. However, as shown in Fig. 14 (top, right), the rank of the positional subspace decreases sharply over depth, eventually collapsing to a 2D plane in the final layers. This compression is reminiscent of a transition from place cell-like coding schemes O’Keefe & Dostrovsky (1971), where individual units encode specific spatial locations, to axis coding Chang & Tsao (2017) where each dimension encodes the continuous value of horizontal or vertical coordinates. This compression is also visible in the PCA projection of positional embeddings themselves (Figure˜15) that eventually collapse to a smooth 2D sheet in the final layers.
@@ 7.1 Theoretical account
Prior work provides evidence that situates MRH within a convex-analytic tradition, which we organize into three strands. First (i), in piecewise-linear networks, ReLU-based architectures partition input space into convex polyhedral linear regions (often unbounded). Theoretical analyses of this partition structure have established upper and lower bounds (and in some cases exact counts) on the number of linear regions and have clarified depth–width trade-offs; in parallel, the spline perspective makes the underlying affine-template structure explicit Montúfar et al. (2014); Telgarsky (2015); Serra et al. (2018); Raghu et al. (2017); Balestriero et al. (2018); Balestriero & Baraniuk (2020). Empirically, prior studies have found that trained networks realize far fewer regions than the maximal theoretical counts and that the realized tessellations depend sensitively on optimization, yielding structured yet parsimonious decompositions Hanin & Rolnick (2019); Zhang & Wu (2020); related interpretability work have exploited this polyhedral structure by enumerating regions to extract exact piecewise-linear rules Black et al. (2022); Chu et al. (2018). Second (ii), in representation space, recent analyses demonstrate convex organization of activations and architecture-specific convex projections Tvetkova et al. (2025). This observation dovetails with results in population geometry indicating that network activity concentrates on low-dimensional manifolds with structured variability Chung (2021); Cohen et al. (2020). Relatedly, recents concepts extraction methods such as k-Deep Simplex Tankala et al. (2023) and SpaDE Hindupur et al. (2025) constrain hidden representations to lie in the convex hull of a learned dictionary (i.e., as convex combinations of simplex vertices), thereby reinforcing a prototype centric view of representation geometry. Third (iii), in language models, recent work has shown that categorical and hierarchical concepts admit polytopal encodings whose geometric relations mirror semantic relations Park et al. (2025).

### v2

@@ 1 Introduction
Trained in a self-supervised manner on a massive corpus of unlabeled images, DINOv2 has become a showcase of emergent visual representations Oquab et al. (2023); Darcet et al. (2023). Its embeddings support diverse downstream tasks: fine-grained classification Chiu et al. (2024), segmentation Liu et al. (2023), monocular depth estimation Mao et al. (2024); Cui et al. (2024), object tracking Faber et al. (2024); Tumanyan et al. (2024), even robotic perception Kim et al. (2024). They have also proven valuable in non-discriminative settings, acting as priors for generative models Yu et al. (2025), proxies for distributional similarity Stein et al. (2023), tools for revealing blindspots Bohacek et al. (2025). The embeddings are not just visually rich, but robust: they power video world models that predict physical dynamics from frames Baldassarre et al. (2025), and can be aligned with language to achieve open-vocabulary segmentation and zero-shot classification despite having never been trained on text Jose et al. (2025). DINOv2 also transfers robustly across domains Zhang et al. (2023); Zhou et al. (2024), including satellite imagery Waldmann et al. (2025) and medical scans Ayzenberg et al. (2024), underscoring that the model has internalized a deeply structured representation of the visual world. Yet despite these successes, the nature of DINOv2’s internal representations remains unclear. What, precisely, is encoded? Which features are available Hermann et al. (2023) to downstream tasks, and how are they geometrically organized? More fundamentally, how can a label-free objective give rise to such effective few-shot readout?
@@
To address these questions, we draw on vision explainability, a field that has developed both an empirical toolkit and theoretical accounts for probing large vision models Doshi-Velez and Kim (2017); Gilpin et al. (2018); Fel (2024). Early work emphasized attribution-based methods Simonyan et al. (2013); Zeiler and Fergus (2014); Bach et al. (2015); Springenberg et al. (2014); Smilkov et al. (2017); Sundararajan et al. (2017); Selvaraju et al. (2017); Fong and Vedaldi (2017); Fel et al. (2021); Novello et al. (2022); Muzellec et al. (2024), which visualize “where” a model looks or how sensitive it is to local perturbations. While often compelling, such techniques provide only a surface-level view: they do not reveal “what” internal features are being computed, and their usefulness is limited Hase and Bansal (2020); Hsieh et al. (2021); Nguyen et al. (2021); Colin et al. (2021); Kim et al. (2022); Sixt et al. (2020). To move beyond attribution, the community turned to concept-based explanations Kim et al. (2018); Poeta et al. (2023), which aim to extract interpretable latent dimensions from a model’s representation space Bau et al. (2017); Ghorbani et al. (2019); Zhang et al. (2021); Fel et al. (2023c); Graziani et al. (2023); Vielhaben et al. (2023); Kowal et al. (2024a; b); Fel et al. (2023b). A natural first attempt was to identify individual neurons as carriers of concepts. However, this view quickly ran into challenges: neurons are polysemantic, depend on arbitrary bases, and implicitly assume the number of features is bounded by the number of units Arora et al. (2018); Elhage et al. (2022); Gorton (2024). These limitations motivated methods that identify concepts as distributed features, without requiring alignment to single neurons Ghorbani et al. (2017); Zhang et al. (2021); Fel et al. (2023c). A first theoretical account soon followed in the form of phenomenology by Elhage et al. (2022): in a $d$-dimensional space, neural networks can encode exponentially many nearly orthogonal features by representing each as a sparse combination of neurons. This phenomenology crystallized into the Linear Representation Hypothesis (LRH), which holds that models contain many more features than neurons, arranged as sparse, quasi-orthogonal directions Park et al. (2024). If the LRH is valid, then concept extraction amounts to an overcomplete dictionary learning problem Fel et al. (2023b). In this framework, the activation space of a layer is factorized into a dictionary basis and sparse codes. Sparse autoencoders (SAEs) Makhzani and Frey (2014); Cunningham et al. (2023); Bricken et al. (2023); Joseph et al. (2025a; b); Klindt et al. (2025) are one practical instantiation of this idea, enforcing sparsity through a simple encoder-decoder architecture. When applied to pretrained models such as DINOv2, SAEs uncover a rich library of patterns, which we refer to as concepts.
@@ 2 Linear Representation Hypothesis and Operationalization
A recurring phenomenology of large models is that their representational capacity vastly exceeds the number of neurons: in a $d$-dimensional space, they can encode exponentially many features by representing each as a sparse linear superposition Arora et al. (2018); Elhage et al. (2022). Empirically, such features behave as nearly orthogonal directions Papyan et al. (2020), active only in restricted contexts, while neurons themselves are polysemantic Nguyen et al. (2016). This motivates two conclusions: (i) neurons are not the appropriate locus of interpretability, and (ii) one must recover the latent basis along which the model effectively operates. Beyond empirical characterization, such geometry can be motivated by compression: when features are arranged with minimal coherence, it maximizes the number of retrievable features while minimizing destructive interference. This principle extends beyond neural activations to fundamental problems in discrete geometry and optimization theory. It connects to classical sphere packing problems: Tammes’s problem Mooers (1994), which seeks optimal angular separation of points on a sphere (originally motivated by pollen grain morphology, where surface protrusions must be optimally spaced for aerodynamic dispersal efficiency); Thomson’s problem Thomson (1904); Bowick et al. (2002), which minimizes Coulomb electrostatic repulsion energy between charged particles constrained to a sphere; and the spherical code problem Delsarte et al. (1991), which maximizes minimum distance between codewords on the unit sphere for error correction. These problems share a common mathematical substrate: the geometric structure formalized as Grassmannian frames in signal processing theory Strohmer and Heath Jr (2003). This object underlies the optimization principle of minimizing mutual coherence (the maximum absolute inner product between distinct normalized vectors). These mathematical convergences suggest that neural networks may naturally approximate such optimal geometric configurations to (i) maximize representational capacity while (ii) minimizing cross-feature interference. This geometric phenomenology crystallizes into the Linear Representation Hypothesis Elhage et al. (2022); Costa et al. (2025) that we recall here:
@@
A persistent challenge, however, is stability: naïve SAEs produce inconsistent features across runs, undermining interpretability Paulo and Belrose (2025); Papadimitriou et al. (2025). To address this, we adopt a stable SAEs Fel et al. (2025), which constrain each dictionary atom to lie in the convex hull of real activations. This guarantees that atoms remain in-distribution and yields reproducible, geometrically faithful dictionaries. Formally, let $(\mathcal{X},\mathcal{F},\mathbb{P})$ denote the probability space of natural images, $\mathcal{X}\subset\mathbb{R}^{H\times W\times 3}$. For a pretrained Vision Transformer $\bm{f}:\mathcal{X}\to\mathbb{R}^{t\times d}$, any image $\bm{x}\sim\mathbb{P}$ yields activations $\bm{a}=\bm{f}(\bm{x})\in\mathbb{R}^{t\times d}$, i.e. $t$ token embeddings of dimension $d$. For a batch $\bm{X}=(\bm{x}_{i})_{i=1}^{n}$, concatenating all tokens gives $\bm{A}\in\mathbb{R}^{nt\times d}$. Our objective is to factorize $\bm{A}$ into sparse codes $\bm{Z}\in\mathbb{R}^{nt\times c}$ and a dictionary $\bm{D}\in\mathbb{R}^{c\times d}$ with
@@ 5 Statistics and Geometry of Concepts
However, superimposed on this sparse backdrop are three outlier concepts that exhibit an anomalously dense activation pattern Sun et al. (2025), firing across the entire dataset. Upon inspection, these concepts are not arbitrary artifacts, but instead they encode positional information: specifically indicating whether a token lies on the left, right, or bottom of the input image grid. Their ubiquity reflects the persistent relevance of spatial structure within the model, and their emergence aligns with our decoding experiments, which reveal that positional information remains linearly decodable up to the intermediate layers (see Appendix˜I) and form a 2-dimensional subspace at the penultimate layer (see Section˜6). This finding adds nuance to the assumption that the concept space is strictly sparse. Instead, we observe a hybrid representational regime where a small set of universally-active, (here spatially grounded features) coexists with a broader ensemble of highly selective, image-contingent activations, resembling sparse and dense low-rank representations Jiang and Lai (2014); Pramanik et al. (2020).
@@
Having examined the statistics of activations, we now turn to the geometry of the learned dictionary.Figure˜10 summarizes structural properties of the concept basis $\bm{D}$, including pairwise dot products, singular value spectrum, and sparsity scores. Local Coherence. Panel (A) reveals that relative to the isotropic random vectors (e.g. at initialization), the dictionary does not move towards the uniform incoherence of an ideal Grassmannian frame. Instead, the SAE dictionary moves to the other direction and exhibits higher coherence: while most atoms are weakly correlated, a non-negligible fraction form tight clusters with stronger alignment. This geometry departs from the assumptions of the LRH. Rather than approximating a Johnson–Lindenstrauss like embedding Larsen and Nelson (2017), [^4] The Johnson–Lindenstrauss (JL) lemma Johnson et al. (1984) implies that $p=\exp(\mathcal{O}(m\varepsilon^{2}))$ vectors can be embedded into $\mathbb{R}^{m}$ such that all pairwise inner products are preserved up to a small distortion $\varepsilon>0$. This supports the feasibility of near-orthogonality even in regimes where $p\gg m$. the model appears to favor structured redundancy, potentially to support specialization or compositional reuse. Moreover, a pattern of interest emerges upon inspecting the tail of the dot-product distribution: a small but significant number of concept pairs are nearly antipodal, i.e., $\bm{D}_{i}\approx-\bm{D}_{j}$ (see Figure˜11, A). This go against the assumption of a Grassmannian-like dictionary  Strohmer and Heath Jr (2003). Instead, it suggests that DINO sometimes exploits paired directions to encode semantically opposed features such as “left vs right” or “white vs black” along shared lines but with opposite signs. Practically, such antipodal structure should be accounted when using cosine similarity in downstream analyses, particularly in data filtering, clustering, or retrieval contexts. Global Anisotropy. The singular value spectrum in Panel (B) shows a sharp decay, indicating that the dictionary spans a low-dimensional subspace despite its overcomplete size. This again suggests an anisotropic allocation of representational capacity: some directions are densely populated, while others are sparsely or not at all represented. One potential explanation is that the activations $\bm{A}$ themselves have anisotropic covariance, thus to reconstruct them $\bm{A}\approx\bm{Z}\bm{D}$ with L2 error, more atoms are allocated to the more important and higher variance subspace. [^5] A speculative hypothesis is that the affine components of the last LayerNorm deforms the activation space into an ellipsoidal geometry, locally increasing curvature along some directions. In such regions, dictionary learning may require more atoms to adequately cover the manifold with low reconstruction error. While this remains to be formalized, similar intuitions appear in manifold tiling and sparse coding contexts Peyré (2009). Distributed Encoding. Finally, Panels (C) and (D) show that the dictionary atoms are not aligned to individual units: their Hoyer scores lie far below the one-hot bound. Compared to random baselines, SAE atoms are slightly sparser but remain broadly distributed—supporting the view that concept representations are shared across the population rather than axis-aligned Elhage et al. (2022); Colin et al. (2024).
@@ 6 The Shape of an Image
A natural hypothesis is that the local connectedness observed in PCA maps, illustrated in Figure˜13 and Figure˜12, is inherited from positional encodings. After all, spatial position is the only structured signal available at initialization, and remains entangled with token identity throughout the network. To test this, we begin by analyzing the evolution of positional information across layers. We trained linear decoders to decode the token position from each layer’s activations and extracted the corresponding decoding vectors at each spatial location $(i,j)$ to form a position encoding matrix $\bm{P}\in\mathbb{R}^{256\times d}$ at each layer (see Appendix˜I for more details). In early layers, this positional embedding matrix $\bm{P}$ is effectively high-rank: we can linearly decode the precise $(i,j)$ location of a token with near-perfect accuracy (see Fig. 14, Top). This aligns with the model’s need for precise positional resolution during early attention operations. However, as shown in Fig. 14 (top, right), the rank of the positional subspace decreases sharply over depth, eventually collapsing to a 2D plane in the final layers. This compression is reminiscent of a transition from place cell-like coding schemes O’Keefe and Dostrovsky (1971), where individual units encode specific spatial locations, to axis coding Chang and Tsao (2017) where each dimension encodes the continuous value of horizontal or vertical coordinates. This compression is also visible in the PCA projection of positional embeddings themselves (Figure˜15) that eventually collapse to a smooth 2D sheet in the final layers.
@@ 7.1 Theoretical account
Prior work provides evidence that situates MRH within a convex-analytic tradition, which we organize into three strands. First (i), in piecewise-linear networks, ReLU-based architectures partition input space into convex polyhedral linear regions (often unbounded). Theoretical analyses of this partition structure have established upper and lower bounds (and in some cases exact counts) on the number of linear regions and have clarified depth–width trade-offs; in parallel, the spline perspective makes the underlying affine-template structure explicit Montúfar et al. (2014); Telgarsky (2015); Serra et al. (2018); Raghu et al. (2017); Balestriero and others (2018); Balestriero and Baraniuk (2020). Empirically, prior studies have found that trained networks realize far fewer regions than the maximal theoretical counts and that the realized tessellations depend sensitively on optimization, yielding structured yet parsimonious decompositions Hanin and Rolnick (2019); Zhang and Wu (2020); related interpretability work have exploited this polyhedral structure by enumerating regions to extract exact piecewise-linear rules Black et al. (2022); Chu et al. (2018). Second (ii), in representation space, recent analyses demonstrate convex organization of activations and architecture-specific convex projections Tvetkova et al. (2025). This observation dovetails with results in population geometry indicating that network activity concentrates on low-dimensional manifolds with structured variability Chung (2021); Cohen et al. (2020). Relatedly, recents concepts extraction methods such as k-Deep Simplex Tankala et al. (2023), MFAs Shafran et al. (2026) and SpaDE Hindupur et al. (2025) constrain hidden representations to lie in the convex hull of a learned dictionary (i.e., as convex combinations of simplex vertices), thereby reinforcing a prototype centric view of representation geometry. Third (iii), in language models, recent work has shown that categorical and hierarchical concepts admit polytopal encodings whose geometric relations mirror semantic relations Park et al. (2025; 2026).